OpenStudy (anonymous):

Find y" in terms of x and y by implicit differentiation (x^3)+(y^3)=4^3 So for this is what I have the first derivative will be (3x^2)+ (3y^2)(dy/dx). I know that before I take the second derivative I need to solve for y'. But, I don't know how to do this Please Help.

OpenStudy (espex):

Solve the equation for dy/dx. \[3y^2\frac{dy}{dx} = -3x^2\]

OpenStudy (anonymous):

(dy/dx) = (-x^2/y^2)

OpenStudy (espex):

Now you can substitute back and take your second derivative.

OpenStudy (anonymous):

so then you would get it down to -3x^4/y^2 - 3y^2x^2/y^2 correct?

OpenStudy (espex):

Substitute what you found for dy/dx into the derivative \[3x^2 + 3y^2\frac{dy}{dx} = 0 \rightarrow 3x^2 + 3y^2\frac{x^2}{y^2} = 0\]

OpenStudy (espex):

Correction: That should have been \[-\frac{x^2}{y^2}\] for dy/dx